# Review of Demographic Bias in Face Recognition

### v1

@@ IV Assessment of Demographic Bias in FR
For an FR system, the FMR represents the proportion of matching scores where a non-mated pair of biometric samples (i.e. samples from different subjects) is incorrectly classified as a match, while FNMR corresponds to the scenarios where a mated pair (i.e. samples of the same subject) is incorrectly classified as non-matching. These metrics are essential for evaluating the performance of FR algorithms– where lower values of error rates are preferred. These errors are determined by the score threshold (often denoted as $\tau$) that binarizes the similarity score into match or no-match decision. As illustrated in Figs. 1(a)1(c), the overlap between distributions of mated scores (shown in green) and non-mated scores (shown in red) defines the regions where classification errors (i.e. false matches and false non-matches) occur.
@@
The choice of score threshold plays significant role in FMR and FNMR. A higher threshold reduces FMR by limiting the likelihood of imposter pairs being classified as matches but increases FNMR by rejecting more genuine pairs, and vice-versa. However, it is essential to distinguish between score overlap and binary decision-making. The overlap reflects the inherent ambiguity in the score distributions, while binarized decisions result from applying the threshold to assign matches or non-matches. For instance, two demographic groups with identical overlap may exhibit different FMR and FNMR due to variations in their score distribution shapes or population characteristics. Similarly, the distributions (mated, non-mated, or both) for two demographic groups can be significantly different albeit exhibiting same error rates.
The demographic bias in FR system leads to variations in score distributions and their overlaps across different demographic groups. Such disparities inherently result in different FMR and FNMR values for each group when a single (global) threshold is used. Fig. 3 illustrates score distributions of different groups may differ, leading to unequal error rates. In this section we briefly review performance measures designed for demographic-aware assessments of bias in FR systems. Given the overlap and distinct characteristics of scores and decisions, establishing a well-defined evaluation framework is crucial for accurate analysis. Howard et al.  [6] introduced the concepts of differential performance and differential outcomes, providing two key terms that aid in achieving a precise understanding and categorization of assessment metrics.
@@
Quantifying demographic bias, in terms of both- demographic performance and outcomes, is critical for developing fair and reliable FR systems. By analyzing the overlap of distributions and error disparities, one can identify specific areas requiring intervention to ensure equity across diverse user groups.
@@
Since the 2019 Face Recognition Vendor Test (FRVT) [^4] Since 2023, the FRVT initiative has been restructured into the Face Recognition Technology Evaluation (FRTE) and Face Analysis Technology Evaluation (FATE) programs. report, NIST has included demographic effects in FR algorithms [15]. This assessment involved comparing non-mated pairs within the same demographic group, setting thresholds for algorithms to achieve an FMR of 0.001 for white males (since this demographic typically associated with the lowest FMR). The report [15] offered a comprehensive analysis of recognition processes and identified areas where demographic effects might occur. To quantify demographic disparity, NIST initially employed Inequity Ratios (IR), calculating the ratio of maximum to minimum of FMR and FNMR across demographic groups. However, considering potentially large range of the error rates, these ratios can become numerically unstable, especially in extreme cases. To alleviate this shortcoming, NIST has considered few modified versions of Inequity Ratios such as adjusting the score threshold ($\tau$) score, incorporating fixed constants in the denominator, or expressing worst-case error rates relative to arithmetic or geometric means [65]. Using the geometric mean is particularly advantageous due to its extended range over FMR/FNMR values. Another possible approach involves referencing a standard FMR or FNMR in the denominator, which inherently resolves stability issues while providing more robust evaluations of demographic bias.
@@
The Fairness Discrepancy Rate (FDR) is one of the initial efforts of quantifying bias in FR [66]. For an FR system using a single decision threshold, the FDR combines FMR and FNMR using weighted sums, and evaluates fairness through a unified measure that captures the trade-offs between both error rates. The FDR requires two hyper-parameters: one for the score threshold, and another for defining relative importance of FMR over FNMR. Thus, it offers flexibility of assessing the fair nature of the model at pre-defined score threshold and application-dependent weighing of false matches to false non-matches. In a similar vein, to define the balance between false matches and false non-matches, NIST proposed an approach to calculate the Inequity Measure by raising the terms representing demographic disparities to specific exponents (serving as weights) and then multiplying them.
Bias Assessment Metric Ref Publication Year Description
@@
TABLE III: Summary of bias assessment metrics employed in demographic analysis of face recognition systems.
@@
Schuckers et al. highlighted the importance of accounting for statistical variation when evaluating fairness in FR systems [67]. They noted that the differences among demographic groups can arise either from actual performance disparities or by chance due to sampling variability, leading to potential Type-I errors. To address this, they proposed two statistical methodologies: a bootstrap-based hypothesis test and a simpler test methodology tailored for non-statistical audiences. Their study also conducted simulations to explore the relationship between margin of error and factors such as the number of subjects, attempts, correlation between attempts, underlying FNMRs, and the number of demographic groups. In [69], the researchers at MdTF proposed new metric for assessment of bias based on demographic outcomes. Their metric, GARBE (Gini Aggregation Rate for Biometric Equitability), is inspired by the Gini coefficient– which has a long history of use as a dispersion measure in socio-economic context. The GARBE evaluates statistical dispersion in error rates and emphasizes equitable treatment across demographic groups. Similar to the FDR and IR, this metric combines weighted contributions of FMR and FNMR to produce a single fairness score for a given fixed score threshold.
@@
Villalobos et al. proposed the Mean Absolute Percentage Error (MAPE) as a metric to quantify differences in error rates across demographic groups [72]. MAPE measures the relative deviation of FMRs from a policy-defined FMR, ensuring that low error rates for one group do not mask higher error rates for another. High deviations in error rates, particularly towards lower values of FMR, can negatively impact the system by increasing FNMR. A MAPE score of zero indicates that all demographic groups achieve the desired FMR, making it an effective metric for fairness evaluation. The Sum of Group Error Differences (SEDG) was introduced as the fairness assessment metric address disparities in biometric verification systems in [70]. The SEDG calculates relative deviations in FMR and FNMR across demographic groups from the FMR/ FNMR of global scores. They consider the Equal Error Rate (EER) threshold as a reference to compute the error rates. In other words, it adapts a relative difference formula to quantify demographic bias by comparing individual group performances to a global standard. Authors argue that by incorporating both within-demographic (WDI) and cross-demographic (CDI) interactions, SEDG is able to provides provide better understanding of the magnitude and type of bias making it a versatile measure.
Relatively fewer attempts have been made to assess the demographic bias at score-level (ie, based on differential performance). Kotwal and Marcel introduced three fairness evaluation measures that emphasize the separation, compactness, and distribution of genuine and impostor scores [68]. Unlike conventional approaches that depend on system accuracy, these measures focus on assessing differential performance without requiring external parameters such as score thresholds. By examining how well the match is, rather than merely determining a match, this approach provides a more nuanced evaluation of demographic fairness. Additionally, they also discussed a weighted fusion strategy to improve relative contributions from under-represented groups, addressing the challenges posed by imbalanced datasets. Building on the work from [68], Solano et al. developed the Comprehensive Equity Index (CEI)– combining error rate differences and recognition score distribution disparities [71]. The CEI enhances bias quantification methods by considering both the distribution tails and overall shapes of score distributions, enabling the detection of subtle biases across demographic groups. They also conducted experiments on high performing FR systems (as per NIST evaluations) using real challenging datasets. Their experiments showed that CEI was able to effectively capture the demographic bias on several challenging datasets with several covariates.
@@
Table III provides a brief summary of the bias assessment metrics discussed in this section.
Figure 4: Illustration of categories of methods of bias mitigation in FR.
@@ V Bias Mitigation in Face Recognition Systems
Pre-processing methods aim to address bias at the level of training data by modifying or augmenting it to reduce discriminatory patterns or imbalances before the data is used for training. This approach is particularly relevant when representational disparities in the data contribute to demographic bias. The in-processing methods focus on modifications to the FR model during training or fine-tuning phase, often by incorporating constraints or objectives that optimize fairness without significantly compromising recognition accuracy. Finally, post-processing techniques involve adjustments to the output of trained models to ensure fairness across demographic groups. These techniques modify the results, such as classification scores or decision thresholds, to achieve equitable performance without altering the underlying model. Fig. 4 summarizes the categorization of bias mitigation methods– that correlates with the corresponding stages of FR pipeline.
@@ V-A Pre-Processing Methods
Pre-processing methods, also known as data-based methods, focus on modifying the biometric samples before feeding into the FR system. These techniques aim to normalize the characteristics of the data to make them robust for subsequent feature extraction. Most of the methods in this category can also be regarded as special case of data augmentation, specifically designed to reduce demographic biases in the training data.
In [23], Klare et al. showed that FR algorithms were performing worse for female, Black, and younger individuals. To address this concern, they proposed two mitigation strategies based on selection of training data. First, training models on specific demographic cohorts to enhance recognition for those groups, and second, implementing a dynamic face matcher selection approach, where different algorithms, each trained on distinct demographic groups, would be chosen based on probe information. A similar approach was followed by Deb et al. in the context of longitudinal study of FR [81]. Using the data of more than 900 children captured over time, they demonstrated that recognition accuracy decreases as the time gap between image captures grows. To address this, they fine-tuned FaceNet [82] on a separate child face dataset, showing improved accuracy. The authors advocated the need for tailored training and evaluation of FR systems for different age groups. Lu et al. [32] investigated the influence of covariates—such as age, gender, pose, and skin tone— toward the performance of face verification. Their findings showed that using gender information to curate training data improves performance, particularly at low FMRs. In [83], Kortylewski et al. demonstrated the effectiveness of using synthetic face images to mitigate the bias arising from real-world datasets. By utilizing synthetic data for pre-training the FR model, they showed that the negative effects of dataset bias, particularly with regard to pose variations, could be significantly reduced. Their experiments revealed that pre-training the CNNs with synthetic data, the need for real-world data can be reduced by up to 75%. Although this study did not directly address demographic bias, we include it in the review as it highlights the potential of synthetic data in improving the generalization performance of FR systems and reducing (non-demographic) dataset bias.

### v2

@@ IV Assessment of Demographic Fairness in FR
For an FR system, the FMR represents the proportion of matching scores where a non-mated pair of biometric samples (i.e., samples from different subjects) is incorrectly classified as a match, while FNMR corresponds to the scenarios where a mated pair (i.e., samples of the same subject) is incorrectly classified as non-matching. These metrics are essential for evaluating the performance of FR algorithms– where lower values of error rates are preferred. These errors are determined by the score threshold (often denoted as $\tau$) that binarizes the similarity score into match or no-match decision. As illustrated in Figs. 1(a)1(c), the overlap between distributions of mated scores (shown in green) and non-mated scores (shown in red) defines the regions where classification errors (i.e., false matches and false non-matches) occur.
@@
The choice of score threshold plays a significant role in FMR and FNMR. A higher threshold reduces FMR by limiting the likelihood of imposter pairs being classified as matches but increases FNMR by rejecting more genuine pairs, and vice-versa. However, it is essential to distinguish between score overlap and binary decision-making. The overlap reflects the inherent ambiguity in the score distributions, while binarized decisions result from applying the threshold to assign matches or non-matches. For instance, two demographic groups with identical overlap may exhibit different FMR and FNMR due to variations in their score distribution shapes or population characteristics. Similarly, the distributions (mated, non-mated, or both) for two demographic groups can be significantly different albeit exhibiting same error rates.
The demographic bias in FR system leads to variations in score distributions and their overlaps across different demographic groups. Such disparities inherently result in different FMR and FNMR values for each group when a single (global) threshold is used. Fig. 3 illustrates score distributions of different groups may differ, leading to unequal error rates. In this section we briefly review performance measures designed for demographic-aware assessments of fairness in FR systems. Given the overlap and distinct characteristics of scores and decisions, establishing a well-defined evaluation framework is crucial for accurate analysis. Howard et al.  [6] introduced the concepts of differential performance and differential outcomes, providing two key terms that aid in achieving a precise understanding and categorization of assessment metrics.
@@
Quantifying demographic bias or fairness, in terms of both demographic performance and outcomes, is critical for developing fair and reliable FR systems. By analyzing the overlap of distributions and error disparities, one can identify specific areas requiring intervention to ensure equity across diverse user groups.
@@
Since the 2019 Face Recognition Vendor Test (FRVT) [^2] Since 2023, the FRVT initiative has been restructured into the Face Recognition Technology Evaluation (FRTE) and Face Analysis Technology Evaluation (FATE) programs. report, NIST has included demographic effects in FR algorithms [15]. This assessment involved comparing non-mated pairs within the same demographic group, setting thresholds for algorithms to achieve an FMR of 0.0001 for white males (since this demographic typically associated with the lowest FMR). The report [15] offered a comprehensive analysis of recognition processes and identified areas where demographic effects might occur. To quantify demographic disparity, NIST initially employed Inequity Ratios (IR), calculating the ratio of maximum to minimum of FMR and FNMR, at score threshold $\tau$, across demographic groups as shown in Eq. 1.
$\text{IR}(\tau)=\left(\frac{\max_{d_{i}}\text{FMR}_{d_{i}}(\tau)}{\min_{d_{j}}\text{FMR}_{d_{j}}(\tau)}\right)^{\alpha}\times\left(\frac{\max_{d_{i}}\text{FNMR}_{d_{i}}(\tau)}{\min_{d_{j}}\text{FNMR}_{d_{j}}(\tau)}\right)^{\beta},$ (1)
–where, the subscripts $d_{i},d_{j}$ refer to the demographic group; and $\alpha,\beta$ are weighing coefficients.
However, considering potentially large range of the error rates, these ratios can become numerically unstable, especially in extreme cases. To alleviate this shortcoming, NIST has considered few modified versions of Inequity Ratios such as adjusting the score threshold ($\tau$), incorporating fixed constants in the denominator, or expressing worst-case error rates relative to arithmetic or geometric means [85]. Using the geometric mean is particularly advantageous due to its extended range over FMR/FNMR values. Another possible approach involves referencing a standard FMR or FNMR in the denominator, which inherently resolves stability issues while providing more robust evaluations of demographic fairness.
@@
The Fairness Discrepancy Rate (FDR) is one of the initial efforts of quantifying fairness in FR [86]. For an FR system using a single decision threshold, the FDR combines FMR and FNMR using weighted sums, and evaluates fairness through a unified measure that captures the trade-offs between both error rates. As described by Eq. 2, the FDR requires two hyper-parameters: one for the score threshold ($\tau$), and another for defining relative importance ($\alpha$) of FMR over FNMR. Thus, it offers flexibility of assessing the fair nature of the model at pre-defined score threshold and application-dependent weighing of false matches to false non-matches.
$A(\tau)=\max\left|\text{FMR}_{d_{i}}(\tau)-\text{FMR}_{d_{j}}(\tau)\right|,$
$B(\tau)=\max\left|\text{FNMR}_{d_{i}}(\tau)-\text{FNMR}_{d_{j}}(\tau)\right|,$
$\text{FDR}(\tau)=1-\left(\alpha A(\tau)+(1-\alpha)B(\tau)\right).$ (2)
In a similar vein, to define the balance between false matches and false non-matches, NIST proposed an approach to calculate the Inequity Measure by raising the terms representing demographic disparities to specific exponents (serving as weights) and then multiplying them.
TABLE IV: Summary of fairness assessment metrics employed in demographic analysis of face recognition systems.
Assessment Metric Ref Publication Year Description
@@
ROC Uncertainty Conti & Clémençon [88] 2024 Utilizes normalized uncertainty of the ROC curves to assess fairness.
@@
d-prime ($d^{\prime}$) - - Quantifies the separation between score distributions of different demographic groups; higher values indicate greater separability.
@@
Schuckers et al. highlighted the importance of accounting for statistical variation when evaluating fairness in FR systems [87]. They noted that the differences among demographic groups can arise either from actual performance disparities or by chance due to sampling variability, leading to potential Type-I errors. To address this, they proposed two statistical methodologies: a bootstrap-based hypothesis test and a simpler test methodology tailored for non-statistical audiences. Their study also conducted simulations to explore the relationship between margin of error and factors such as the number of subjects, attempts, correlation between attempts, underlying FNMRs, and the number of demographic groups. In [88], the authors propose a statistical framework for assessing fairness using Receiver Operating Characteristics (ROC) curves and associated uncertainty measures derived from U-statistics. They argue that the ROC curve is not only central to evaluating system performance but also serves as a key tool for analyzing fairness in similarity scoring. They established asymptotic guarantees for empirical ROC curves and fairness-relevant metrics, emphasizing the importance of accurately quantifying uncertain. Notably, they highlight that conventional bootstrap methods may misrepresent uncertainty due to the U-statistical nature of FMR and FNMR, and recommend a recentering technique to ensure valid inference. In [90], the researchers at MdTF proposed new metric for assessment of bias based on demographic outcomes. Their metric, GARBE (Gini Aggregation Rate for Biometric Equitability), is inspired by the Gini coefficient– which has a long history of use as a dispersion measure in socio-economic context. The GARBE evaluates statistical dispersion in error rates and emphasizes equitable treatment across demographic groups. Similar to the FDR and IR, this metric combines weighted contributions of FMR and FNMR to produce a single fairness score for a given fixed score threshold. The formula for GABRE and its constituent Gini coefficient are provided in Eq. 3.
GARBE $=\alpha\,G_{\text{FMR}}(\tau)+(1-\alpha)\,G_{\text{FNMR}}(\tau),\quad\text{where},$
$G_{x}$ $=\left(\frac{n}{n-1}\right)\left(\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}|x_{i}-x_{j}|}{2n^{2}\bar{x}}\right).$ (3)
@@
Villalobos et al. proposed the Mean Absolute Percentage Error (MAPE) as a metric to quantify differences in error rates across demographic groups [93]. MAPE measures the relative deviation of FMRs from a policy-defined FMR, ensuring that low error rates for one group do not mask higher error rates for another. High deviations in error rates, particularly towards lower values of FMR, can negatively impact the system by increasing FNMR. A MAPE score of zero indicates that all demographic groups achieve the desired FMR, making it an effective metric for fairness evaluation. If $\text{FMR}_{p}$ indicates policy-defined FMR, then MAPE for a system using $N$ groups can be computed as:
$\text{MAPE}=\frac{100}{N}\,\sum_{d=1}^{N}\left|\frac{\text{FMR}_{d}-\text{FMR}_{p}}{\text{FMR}_{p}}\right|.$ (4)
The Sum of Group Error Differences (SEDG) was introduced as the fairness assessment metric address disparities in biometric verification systems in [91]. The SEDG calculates relative deviations in FMR and FNMR across demographic groups from the FMR/ FNMR of global scores. They consider the Equal Error Rate (EER) threshold as a reference to compute the error rates. In other words, it adapts a relative difference formula to quantify demographic fairness by comparing individual group performances to a global standard. Authors argue that by incorporating both within-demographic (WDI) and cross-demographic (CDI) interactions, SEDG is able to provides provide better understanding of the magnitude and type of bias making it a versatile measure. For a setup comprising $\mathcal{D}$ set of demographic groups, the SEDG can be obtained as:
$\delta X_{d}=\left|1-\frac{X_{d}(\tau)}{X_{\text{global}}(\tau)}\right|,\cdots\quad X\in(\text{FMR},\text{FNMR}),$
$\text{SED}_{\rm G}=\{\delta\text{FMR}_{d}+\delta\text{FNMR}_{d}\,\,|\,d\in\mathcal{D}\}.$ (5)
Relatively fewer attempts have been made to assess the demographic fairness at score-level (i.e., based on differential performance). Among the various metrics used to quantify fairness in biometric systems, d-prime ($d^{\prime}$) has seen increasing adoption [45, 53, 42]. Originally from signal detection theory, it measures the distance between score distributions of mated and non-mated pairs, and can be adapted to assess separability between demographic group distributions. It is threshold-independent, making it robust in cases where system decision thresholds are not aligned across deployments. If $\mu,\sigma$ represent the mean and standard deviations of the scores, the $d^{\prime}$ is obtained as:
$d^{\prime}=\frac{|\mu_{m}-\mu_{nm}|}{\sqrt{\frac{1}{2}(\sigma_{m}^{2}+\sigma_{nm}^{2})}},$ (6)
—where the subscripts $m$ and $nm$ refer to the distributions of mated and non-mated pairs, respectively.
In [89], Kotwal and Marcel introduced three fairness evaluation measures that emphasize the separation, compactness, and distribution of genuine and impostor scores. Unlike conventional approaches that depend on system accuracy, these measures focus on assessing differential performance without requiring external parameters such as score thresholds. By examining how well the match is, rather than merely determining a match, this approach provides a more nuanced evaluation of demographic fairness. The Separation Fairness Index (SFI) measures the consistency in separation between the mean genuine and impostor scores across groups—higher consistency implies better fairness. The Compactness Fairness Index (CFI) assesses the similarity in intra-group score variability by comparing the standard deviations of genuine and impostor scores across demographics. The Distribution Fairness Index (DFI) evaluates how closely the entire score distribution of each group matches a reference distribution (e.g., the population average), using a statistical divergence measure. Their normal variants are summarized in Eq. 7.
$S(d)=|\mu_{m}(d)-\mu_{nm}(d)|;\text{and }C(d)=\sigma_{m}(d)+\sigma_{nm}(d);$
$\text{SFI}=1-\frac{2}{N}\sum_{d=1}^{N}\left|S(d)-\bar{S}\right|;\quad\text{CFI}=1-\frac{2}{N}\sum_{d=1}^{N}|C(d)-\bar{C}|$
$\text{DFI}=1-\frac{1}{N\,\log_{2}N}\sum_{d=1}^{N}\mathcal{D}_{\rm KL}\left(\mathbb{P}(d)\,||\,\mathbb{P}_{\text{ref}}\right)$ (7)
Here, $N$ denotes the number of demographic groups; $\mu_{m}(d)$ and $\mu_{nm}(d)$ are the mean mated and non-mated scores for group $d$; $\sigma_{m}(d)$ and $\sigma_{nm}(d)$ are their standard deviations; $\mathbb{P}(d)$ is the score distribution for group $d$; $\mathbb{P}_{\text{ref}}$ is the reference distribution (often the average across groups); and $\mathcal{D}_{\rm KL}$ represents the KL divergence. Building on the work from [89], Solano et al. developed the Comprehensive Equity Index (CEI)– combining error rate differences and recognition score distribution disparities [92]. The CEI enhances bias quantification methods by considering both the distribution tails and overall shapes of score distributions, enabling the detection of subtle biases across demographic groups. They also conducted experiments on high performing FR systems (as per NIST evaluations) using challenging, real-world datasets. Their experiments showed that CEI was able to effectively capture the demographic bias on several challenging datasets with several covariates. In compact manner, the CEI can be described as: $\text{CEI}=w_{1}\cdot\text{Tail Disparity}+w_{2}\cdot\text{Central Disparity}$. The disparity calculation is similar to that of DFI expressed in Eq. 7.
@@
Table IV provides a brief summary of the fairness assessment metrics discussed in this section. While several metrics listed here capture similar goals: quantifying inter-group performance disparities, they differ in sensitivity, interpretability, and need for threshold selection. Some, that employ differential FMR/FNMR, are easier to relate to operational performance but may not generalize across systems with differing threshold settings. Others, like d-prime and fairness index measures, focus on distributional shifts and are less sensitive to specific thresholds. As accuracy differentials become narrower with improvements in model architectures and training data, future fairness evaluations should include both threshold-based and threshold-agnostic metrics.
@@ V Bias Mitigation in Face Recognition Systems
Figure 4: Illustration of categories of bias mitigation methods in FR.
@@
Pre-processing methods aim to address fairness at the level of training data by modifying or augmenting it to reduce discriminatory patterns or imbalances before the data is used for training. This approach is particularly relevant when representational disparities in the data contribute to demographic bias. The in-processing methods focus on modifications to the FR model during training or fine-tuning phase, often by incorporating constraints or objectives that optimize fairness without significantly compromising recognition accuracy. Finally, post-processing techniques involve adjustments to the output of trained models to ensure fairness across demographic groups. These techniques modify the results, such as classification scores or decision thresholds, to achieve equitable performance without altering the underlying model. Fig. 4 summarizes the categorization of bias mitigation methods– that correlates with the corresponding stages of FR pipeline. These methods are generally aimed at minimizing accuracy disparities across demographic groups while maintaining overall recognition performance. However, empirical evidence from most studies indicates that efforts to reduce such disparities often lead to a reduction in overall accuracy.
@@ V-A Pre-Processing Methods
Pre-processing methods, also known as data-based methods, focus on modifying the biometric samples before feeding into the FR system. These techniques aim to normalize the characteristics of the data to make them robust for subsequent feature extraction. Most of the methods in this category can also be regarded as special case of data augmentation, specifically designed to reduce demographic disparities in the training data.
In [23], Klare et al. showed that FR algorithms were performing worse for female, Black, and younger individuals. To address this concern, they proposed two mitigation strategies based on selection of training data. First, training models on specific demographic cohorts to enhance recognition for those groups, and second, implementing a dynamic face matcher selection approach, where different algorithms, each trained on distinct demographic groups, would be chosen based on probe information. A similar approach was followed by Deb et al. in the context of longitudinal study of FR [102]. Using the data of more than 900 children captured over time, they demonstrated that recognition accuracy decreases as the time gap between image captures grows. To address this, they fine-tuned FaceNet [103] on a separate child face dataset, showing improved accuracy. The authors advocated the need for tailored training and evaluation of FR systems for different age groups. Lu et al. [34] investigated the influence of covariates–such as age, gender, pose, and skin tone– toward the performance of face verification. Their findings showed that using gender information to curate training data improves performance, particularly at low FMRs. In [104], Kortylewski et al. demonstrated the effectiveness of using synthetic face images to mitigate the accuracy differentials arising from real-world datasets. By utilizing synthetic data for pre-training the FR model, they showed that the negative effects of dataset bias, particularly with regard to pose variations, could be significantly reduced. Their experiments revealed that pre-training the CNNs with synthetic data, the need for real-world data can be reduced by up to 75%. Although this study did not directly address demographic fairness, we include it in the review as it highlights the potential of synthetic data in improving the generalization performance of FR systems and reducing (non-demographic) dataset bias. Yucer et al. [105] proposed a GAN-based adversarial image-to-image augmentation method to generate racially balanced training data at subject level while preserving identity features. Evaluated across multiple loss functions, their approach reduced performance variance across racial groups and improved recognition accuracy for minority subjects.